In a game, $3 is bet on the roll of a 7 sided die. If 3, 4, 5, 6 is rolled, the player wins $2.
a. What is the expected winnings from the game?
b. Is the game fair? Explain why or why not.

Answer by stanbon(75887)

In a game, $3 is bet on the roll of a 7 sided die.
If 3, 4, 5, 6 is rolled, the player wins $2.
a. What is the expected winnings from the game?
Random # values: -3.......2
Probabilities::: 3/7......4/7
---
Expected winnings: (-9 + 8)/7 = -1/7 dollars
-------
b. Is the game fair? Explain why or why not.
Not fair as you can expect to lose 1/7 = $0.1428 cents each time you play.
==================
Cheers,
Stan H.